# Solid-State Dewetting of Polycrystalline Thin Films: a Phase Field Approach

### v1

@@ Abstract
Figure 3b shows a representative morphological evolution during dewetting. It is obtained by a MPF simulation with the idealized initial configuration shown in Fig. 2b. A relatively large $r^{\rm 3D}$ is chosen to ensure film rupture. SSD initiates at junctions, confirming seminal equilibrium considerations [10], other recent simulations [30], and predictions discussed above. Additionally, a detailed account of the dynamics at GB junctions is obtained, accurately capturing morphological evolution, including the formation and growth of non-trivial opening holes. At intermediate stages, as in 2D, grains detach from neighbors and evolve into single-crystalline patches. A similar qualitative behavior, enriched by additional features, emerges at larger aspect ratios; see Fig. 3c. Breakup initiates at triple junctions and rapidly follows at grain boundaries. When the grains detach completely from each other, the dynamics converge to the one expected for a single crystalline patch. A rim forms at the edges (see also Fig. 3a for $r^{\rm 3D}=16$). For the considered aspect ratio, a breakup at the center of the patch is obtained, resembling a classical feature of SSD of single-crystalline patches  [2, 24, 4]. A semi-toroidal shape forms and, driven by curvature gradients, evolves until the central hole eventually closes.
@@
We finally present the case of a polycrystalline thin film of finite size (a patch); see Fig. 4, which reports the SSD dynamics for two different initial conditions. The grains are initialized with varying morphologies and aspect ratios. As observed earlier, rims form at the grain boundaries, while holes open at junctions. Such holes initiate dewetting fronts within the patch, while a continuous rim effectively develops along the patch’s edges, breaking only at later stages. Aside from the rapid disappearance of small grains, the slower dynamics of rims and peripheral GBs, compared to hole opening, are evident along edges with both relatively small and large grains. This can be seen, for instance, by comparing the upper and lower edges in Fig. 4a, and it is also obtained for a more heterogeneous distribution of grains, as shown in Fig. 4b. We thus find that SSD in polycrystalline patches leads to early-stage morphologies characterized by material accumulation between the center and the initial patch boundaries, closely resembling the behavior observed in single crystals [24, 4]. The presence of GBs eventually leads to uncontrolled breaking within the patch for usual polycrystals, where grain contributions are dominated by randomness. However, the design of GB networks may lead to additional degrees of freedom to drive self-assembly in a controlled manner. In this context, we remark that the opening of holes occurs relatively fast compared to the rim retraction mediated by surface diffusion (see, e.g., Fig. 4b).
In conclusion, we demonstrated a MPF tackling the complexity of SSD dynamics in polycrystalline thin films. The simulation results are consistent with independent predictions based on energetic arguments and successfully reproduce the key phenomenology of surface diffusion and curvature-driven GB motion in systems with multiple grains up to equilibrium. We further elucidated key aspects of the onset of dewetting and analyzed the SSD of polycrystalline patches, highlighting the differences and similarities with single-crystalline SSD. Tailoring GB distributions can expand the design space for SSD-driven self-assembly. In this context, our approach offers a versatile tool to explore the associated parameter space. For the sake of generality, we have illustrated the framework using ideal, minimal settings, particularly concerning material parameters. However, by leveraging extended parameterizations and couplings already available in PF models, such as those addressing anisotropic surface or GB energy densities [35, 36, 37] as well as elasticity and plasticity effects [21, 38, 39], the proposed framework can be adapted to material-specific settings.
@@
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Data Availability
The data that support the findings of this article are openly available [40].
@@
Theory on N-gonal Polycrystalline Thin-Films
@@
$\tilde{z}_{n}(\theta,\tilde{\rho})$ $=\sqrt{\sin^{-2}(\theta)-\tilde{\rho}^{2}}-\sqrt{\sin^{-2}(\theta)-\tilde{R}^{2}},$ (S-16)
$\vartheta_{n}(\tilde{\rho})$ =$=\begin{cases}\pi/n&\tilde{\rho}<1\\ \pi/n-\arccos(\tilde{\rho}^{-1})&\tilde{\rho}\geq 1\end{cases}\,.$ (S-17)

### v2

@@ Abstract
Figure 3b illustrates the morphological evolution during dewetting described by MPF simulations, considering an idealized initial configuration as in Fig. 2b and a relatively large $r^{\rm 3D}$ to ensure film rupture. SSD initiates at junctions, confirming seminal equilibrium considerations [10], other recent simulations [30], and quantitative predictions discussed above. Additionally, a detailed account of the dynamics at GB junctions is obtained, including the formation and growth of non-trivial opening holes. At intermediate stages, grains detach from neighbors and evolve into single-crystalline patches. A similar qualitative behavior, enriched by additional features, emerges at larger aspect ratios; see Fig. 3c. Importantly, breakup is found to initiates at triple junctions and rapidly propagates along grain boundaries, forming multiple contact points. When the grains detach completely from each other, the dynamics reduces to single-crystalline SSD. A rim forms at the edges (see also Fig. 3a for $r^{\rm 3D}=16$). For the considered aspect ratio, a breakup at the center of the patch is obtained, resembling a classical feature of single-crystalline SSD [2, 24, 4]. A semi-toroidal shape forms and, driven by curvature gradients, evolves until the central hole eventually closes.
@@
We finally present a novel scenario: SSD of a polycrystalline patch; see Figs. 4a and 4b illustrating the SSD for two patches with varying grain morphologies and aspect ratios. Grooves form at GBs, while holes open at junctions. Such holes initiate dewetting fronts within the patch, while a continuous rim effectively develops along the edges, breaking only at later stages. Aside from the rapid disappearance of small grains, the slower dynamics of rims and peripheral GBs, compared to hole opening, are evident along edges with both relatively small and large grains. We thus find that SSD in polycrystalline patches leads to early-stage morphologies characterized by material accumulation between the center and the initial patch boundaries, closely resembling the behavior observed in single crystals [24, 4], reported in Fig. 4c for completeness. In typical polycrystals, the presence of GBs leads to random, uncontrolled breakup within the patch. However, designing GB networks can provide additional degrees of freedom to control self-assembly. Notably, hole opening occurs much faster than rim retraction driven by surface diffusion (c.f., Fig. 4).
In conclusion, we demonstrated a grandpotential MPF model tackling the geometric complexity of SSD dynamics in polycrystalline thin films. The simulation results are consistent with novel independent predictions based on energetic arguments and characterize the central, non-trivial morphological evolutions. Expressions of the grain critical aspect ratio for dewetting have been derived and represent a theoretical benchmark for scenarios well beyond the simulations reported here. We further elucidated key aspects of the dewetting onset and analyzed the SSD of polycrystalline patches, highlighting the differences and similarities with single-crystalline SSD. Tailoring GB distributions can expand the design space for SSD-driven self-assembly. Our approach provides a versatile tool for exploring the associated parameter space. For the sake of generality, we have illustrated idealized, minimal settings, particularly with respect to material parameters. Nevertheless, such simplified conditions have already proven powerful for understanding and predicting single-crystalline solid-state dewetting [24, 4], and we reported here for the first time the characteristic traits of their polycrystalline counterparts. The considered MPF model still allows for extended parameterizations and couplings compatible PF models, e.g., incorporating anisotropic surface or GB energy densities [39, 40, 41], elasticity, and plasticity effects [21, 42, 43]. Careful assessment against theoretical predictions and experimental observations will be essential to quantitatively assess the influence of these additional contributions. Incorporating specifically shear-coupled GB migration [43] still requires extending the underlying continuum model to fully 3D geometries and accounting for free surfaces of arbitrary morphology. While these remain open challenges, the approach proposed here provides a solid foundation for future developments that fully incorporate capillarity and capture concurrent kinetic processes.
Data Availability
The data supporting the findings of this article are openly available [44].
CRediT authorship contribution statement
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
@@
Theory of N-gonal Polycrystalline Thin-Films
@@
[…]